OpenStudy (anonymous):

Solve: [0,π] sinx-sinxcosx=0

OpenStudy (unklerhaukus):

x=0 is a solution

OpenStudy (anonymous):

your a genius... -.- smh

OpenStudy (anonymous):

cosx=1 x=arcos1

OpenStudy (anonymous):

ohh arc... alright.. how do you do the arc problems again, been ages (review for finals is a feather.. lol)

OpenStudy (unklerhaukus):

\[\sin x-\sin x \cos x=0\]\[\sin x(1-\cos x)=0\] \[\sin x=0\] when \[x\in n\pi\]\[n \in R\], there are only two point in the domain

OpenStudy (anonymous):